The acceleration of a particle is increasing linearly with time t as bt. The particle starts from origin with initial velocity u. The distance travelled by the particle in time t will be:


a) ut+(1/3)bt2


b) ut+(1/2)bt2


c) ut+(1/6)bt3


d) ut+(1/3)bt3


 

Grade:11

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear abhingya

d2s/dt2 =bt

intigrate

V= ds/dt = bt2/2 +c

at t=0 V=U

so U =0 +c

ds/dt = bt2/2 +U

furthet intigrate

s= b/6 t3 +Ut +d

at t=0 s=0

so d=0

S =b/6 t3 +Ut

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Mukunda Nagre
6 years ago
We have to assume acceration =0 when t=0. So we have the equation: a = bt Since a = dv/dt dv/dt = bt Integrating gives: v = (b/2)t² + C When t=0, v = u, so C = u v = (b/2)t² + uSince v = ds/dt ds/dt = (b/2)t² + uIntegrating gives: s = (b/6)t³ + u.t + K When t=0, s=0, so K=0 s= (b/6)t³ + u.t